If you drive to a beach that is 90 miles away, spend 5 hours at the beach and then drive home, and if your average driving speed is 45 miles per hour, what is the complete time of your trip that day?
Answer by mananth(16949)

distance = 90 miles
speed = 45mph
time taken to reach the beach = 90/45 = 2 hours
return time will be 2 hours
time spent on beach = 5 hours
..
total time for the trip = 2+5+2=9 hours